OpenStudy (anonymous):

Megan had some peaches. She gave 4 peaches to her brother. Which expression represents this situation? A. p + 4 B. p – 4 C. p × 4 D. p ÷ 4

OpenStudy (anonymous):

@Crazy_questions

OpenStudy (savannaadams16):

B

OpenStudy (anonymous):

wanna do a nother one?

OpenStudy (anonymous):

another

OpenStudy (savannaadams16):

Sure!

OpenStudy (anonymous):

Which equation means four times a number is equal to 24? A. 24 + 4 = n B. 4 × 24 = n C. 4 + n = 24 D. 4 × n = 24

OpenStudy (savannaadams16):

D

OpenStudy (anonymous):

another one?

OpenStudy (savannaadams16):

Of course.

OpenStudy (anonymous):

What is the value of r? r × 3 × 2 = 3 × 4 × 2 A. 4 B. 3 C. 2 D. 1

OpenStudy (anonymous):

its 2 right?

OpenStudy (anonymous):

guys?

OpenStudy (mathmale):

Hold it, please. This is not a place to exchange answers with no work involved. @savannaadams16 : Please read http://openstudy.com/code-of-conduct. Let p represent the number of peaches that Megan had in the beginning. If she gives away 4 of them, how many does she have left? Write an algebraic expression. You do NOT have enough info to say "2, right?"

OpenStudy (mathmale):

Will Megan have more peaches than before, the same number of peaches, or fewer peaches than before?

OpenStudy (mathmale):

What is the value of r? r × 3 × 2 = 3 × 4 × 2 Simplify the multiplication on both sides. Then SOLVE for r. Show all steps, please.

OpenStudy (anonymous):

i already got that ]:

OpenStudy (mathmale):

Next time, please show all work for those problems with which you want help. Thanks.

OpenStudy (savannaadams16):

I apologize, I didn't know I needed to show my work.

OpenStudy (anonymous):

no need to apologize i just needed to double check :D